OpenStudy (anonymous):

Hello All: I have a team assignment due this evening and I am having the hardest time trying to figure out my problem: Cost of Tuition In 2000, the average cost of tuition and fees at private four-year colleges was $16,200, and in 2005 it was $20,100. Sketch a line that passes through the points (2000, 16200) and (2005, 20100). my question is this: Rate of Change in Tuition Calculate the slope of the line in your graph. Interpret this slope as a rate of change. HELP!!

OpenStudy (anonymous):

take year in x and tution fee in y then initial was (2000,16200) and later was (2005,20100) so our slope will be (20100-16200)/(2005-2000) = 39/5 =7.5 that is rate of fee per year!!!!!!!!

OpenStudy (anonymous):

What is the formula? I have to show my work or how I got the answer.

OpenStudy (anonymous):

This Math 208 is kicking my butt!